OpenStudy (anonymous):

how can i solve exponential equation 50^x = 17

OpenStudy (anonymous):

you stupid boy

OpenStudy (anonymous):

flutter off..

OpenStudy (anonymous):

i need to make the bases equal..

OpenStudy (hitaro9):

So, you're looking to find 50 to what power = 17. Knowing that 17 is less than 50, you can assume that x is going to be less than 1. Now. have you learned about logarithms yet?

OpenStudy (anonymous):

Kind of, its going to be a fraction isnt it

OpenStudy (hitaro9):

You know it's going to be less than 1, so yeah.

OpenStudy (anonymous):

how would solving this be different with equal bases?

OpenStudy (hitaro9):

Okay, so 17 = 50^x can be changed to \[\log_{50}(17)\] = x

OpenStudy (anonymous):

oh i forgot about this!

OpenStudy (anonymous):

so now i solve for x?

OpenStudy (hitaro9):

You need a numeric answer? log50(17) won't do?

OpenStudy (anonymous):

the question is asking: Erika was working on solving the exponential equation 50x = 17; however, she is not quite sure where to start. Using complete sentences, describe to Erika how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (hitaro9):

I mean, I just used the basic definition of logarithms to find the answer. logb(y)=x is the same as b^x = y

OpenStudy (hitaro9):

beyond that I don't really remember how one goes about finding logs.

OpenStudy (hitaro9):

You could say she could just guess and check, but I doubt your teacher wants that.

OpenStudy (anonymous):

i just need to know what the equal base is, i havent sound one yet.

OpenStudy (anonymous):

found

OpenStudy (hitaro9):

If the bases are equal then the problem is really easy If 50^x = 50^2 then x=2 You simply set the exponents equal to one another.

OpenStudy (anonymous):

that does not make any sense to me.. 50^2 is not equal to 17...

OpenStudy (hitaro9):

Well I think it's just a what if kind of thing, unless I'm misunderstanding what they're asking?

OpenStudy (hitaro9):

"describe to Erika how to solve this equation and how solving would be different if the bases were equal." if the bases were equal The bases aren't equal so it's not really relevant to the problem.

OpenStudy (anonymous):

it would look something like this.. (x)^(x)(n)=(x)^(n) i think

OpenStudy (hitaro9):

I'm sorry, I just don't know what the problem is asking. Maybe someone else can help ya'

OpenStudy (anonymous):

ok thanks i think i figured it out 50^x=50^(.34)

OpenStudy (hitaro9):

Alright.